# Waiting around for Unruh

### v1

@@ 1 Introduction
Despite the broad agreement of the circular acceleration effective temperature with the Unruh temperature (1.1), the two differ significantly in $2+1$ spacetime dimensions when the internal energy spacing of the accelerating detector is small and the interaction with the quantum field lasts for a long time [25]. In this regime, the circular acceleration effective temperature is much smaller than the Unruh temperature: for a pointlike two-level quantum detector, in the limit of infinite interaction time and small energy gap, the effective temperature vanishes linearly in the gap [25]. Mathematically, this phenomenon is characteristic of circular motion in $2+1$ dimensions, and it originates in the weak fall-off of the field’s Wightman function along the circular worldline [44]. Physically, this phenomenon is relevant to ongoing experimental efforts to observe the analogue circular motion Unruh effect in superfluid Helium thin-films [45] and Bose-Einstein condensates [24, 27], where the quantum field is effectively $(2+1)$-dimensional.
The purpose of this paper is to investigate to what extent the assumption of an infinitely long interaction duration is responsible for the disparity between the circular motion effective temperature and the Unruh temperature in the small gap regime. In brief: How long does an observer undergoing circular motion in $2+1$ dimensions need to wait to register approximate thermality when the observer’s internal energy spacing is small?
@@
Switching functions with sign changes appear not to have received attention in the literature. However, there is nothing that prevents us a priori from considering such switching functions, or from constructing them in laboratory experiments. In fact, they arise naturally in entanglement harvesting protocols in which causally disconnected local laser pulses become entangled through their interaction with, for example, an electromagnetic field [49] or a Bose-Einstein condensate [50]. In this setting, the switching function is given by the coherent amplitude of the electric field of the laser, which can be given any profile, and in particular, it can be engineered to change sign.
@@ 2.2 Stationary response and the long time limit
Consider now the limit in which the interaction interaction operates for a long time with approximately constant coupling strength. As $\chi$ has by assumption compact support, we may formalise the long time limit by considering a family of switching functions $\chi_{\lambda}$, where $\lambda$ is a positive parameter, such that the support of $\chi_{\lambda}$ is an interval with length proportional to $\lambda$ as $\lambda\to\infty$, and $\chi_{\lambda}$ is approximately constant within this interval as $\lambda\to\infty$, at least in some averaged sense. Specific implementations of this limit were introduced in [46], under the assumption that $\chi_{\lambda}$ is non-negative, and we will employ these implementations in Sections 4 and 6. These implementations satisfy
@@ 2.3 Effective temperature
For uniform linear acceleration with proper acceleration $a$, with the field initially prepared in the Minkowski vacuum, $T(E)$ is independent of the energy $E$ and equal to the Unruh temperature $T_{U}=a/(2\pi)$ [3, 43]. The state $\rho(E)$ (2.8) in this case is thus a genuine thermal Gibbs state in the temperature $T_{U}$, and (2.9) is the detailed balance relation between excitation and de-excitation rates in thermal equilibrium [65, 66]. For other types of non-inertial stationary motion, $\rho(E)$ (2.8) is not a thermal state in the Gibbs sense since $T(E)$ depends $E$; however, for sufficiently narrow intervals in $E$, we may view $\rho(E)$ as an approximate Gibbs state in the energy-dependent temperature $T(E)$. We refer to $T(E)$ as the (infinite time) detailed balance temperature. Studies of $T(E)$ in a range of situations are given in [38, 25, 64, 40, 39, 44, 41].
@@ 3 Asymptotically scaled switching families
using Plancherel’s theorem. The particular case (3.3) follows immediately with $V(E,\omega)=\widehat{W}(E+\omega)$, and the last statement (3.4) is just the fact that the existence of a double limit implies the existence of the limit in $\lambda$ with $E$ held fixed. ∎
@@ Acknowledgements
We thank Benito Juárez-Aubry, Cameron Bunney, Cisco Gooding, Robert Mann, Rick Perche, Bill Unruh and all the members of the Nottingham Gravity Laboratory for stimulating discussions throughout the course of this work. DV-C thanks CONACyT and the WW Smith Trust for financial support. The work of CJF was supported by the Engineering and Physical Sciences Research Council [grant number EP/Y000099/1]. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. The authors have benefited from the activities of COST Action CA23115: Relativistic Quantum Information, funded by COST (European Cooperation in Science and Technology). For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.
@@ B.1 $\alpha>1$
[…]
@@ B.3 $\alpha<1$
$\mathcal{B}(E)=\eta(|E|)\int_{-1}^{1}\mathrm{d}v\,\frac{1-\cos\big(\eta(|E|)v\big{missing})}{\eta^{2}(|E|)v^{2}}|\widehat{\psi}(|E|v)|^{2},$ (B.10)
@@
$\frac{1}{|E|}\int_{-\eta(E)}^{\eta(E)}\mathrm{d}u\,{|\widehat{\chi}(u)|}^{2}$ $=\frac{2S}{|E|}{\textstyle{E_{1}\!\left(\log\bigl(\frac{S}{\eta(E)}\bigr{missing})\right)}}\bigl{(}1+o(1)\bigr{)}$
$=\frac{2S}{|E|}\!\left(\frac{\frac{\eta(E)}{S}}{\log\bigl(\frac{S}{\eta(E)}\bigr{missing})}\right)\bigl{(}1+o(1)\bigr{)}$
$=\frac{2\lambda(E)}{\log(S/|E|)-\log\bigl(\lambda(E)\bigr{missing})}\bigl{(}1+o(1)\bigr{)},$ (C.4)
@@
Finally, combining (D.3a) and (D) gives $I(v)=-\frac{2\sqrt{3}}{\pi^{2}}\log\bigl(1-v\bigr{missing})+O(1)$ as $v\to 1$.
@@
$J(\gamma)=\sqrt{3}\log\bigl(\tfrac{2}{\sqrt{3}}\gamma\bigr{missing})+\int_{0}^{\infty}\mathrm{d}z\left(\frac{1}{\sqrt{1-\operatorname{\text{sinc}}^{2}\!z}}-\frac{\sqrt{3}}{z(1+z)}-1\right)+o(1),$ (D.9)
@@ E.2 $\chi_{\lambda,\delta}$ is an ASSF
$=\frac{1}{2\pi\lambda^{1-\delta}}\int_{-\infty}^{\infty}\mathrm{d}\tilde{\Omega}\,\widehat{f}\bigl{(}\tilde{\Omega}/\lambda^{1-\delta}\bigr{)}\widehat{\chi}\bigl{(}u-\tilde{\Omega}\bigr{)}$, (E.2)
@@ E.3 $\chi_{\lambda,\delta}$ satisfies SFS when $0<\alpha<2/3$ and $\delta>3/2$
$=\int_{{\mathbb{R}}^{2}}\mathrm{d}u\,\mathrm{d}v\,\cos\bigl(\omega(u-v)\bigr{missing})\chi(u/\lambda)\chi(v/\lambda)f(u/\lambda^{\delta})f(v/\lambda^{\delta})$
$=\lambda^{2}\int_{{\mathbb{R}}^{2}}\mathrm{d}x\,\mathrm{d}y\,\cos\bigl(\omega\lambda(x-y)\bigr{missing})\chi(x)\chi(y)f(\lambda^{1-\delta}x)f(\lambda^{1-\delta}y)\,$, (E.6)
where the second equality follows because the contribution from $\sin\bigl(\omega(u-v)\bigr{missing})$ vanishes by antisymmetry under $(u,v)\to(v,u)$, and the third equality follows by the substitution $u=\lambda x$ and $v=\lambda y$. The expression on the left-hand side of (7.8) takes hence the form

### v2

@@ 1 Introduction
In the recent analogue spacetime proposals to observe the circular motion Unruh effect in Bose-Einstein condensates [32, 33] and superfluid Helium thin-films [9], the quantum field that is simulated is a $(2+1)$-dimensional massless scalar field. In this system, the circular acceleration effective Unruh temperature is of the same order of magnitude as the linear acceleration prediction over most of the parameter space, but it is much smaller when the internal energy spacing of the accelerating detector is small and the interaction with the quantum field lasts for a long time [5]. This phenomenon is characteristic of circular motion in $2+1$ dimensions for a massless field, and its mathematical origin is in the weak fall-off of the field’s Wightman function along the circular worldline [54]. The phenomenon hence presents a challenge for experiments that would operate in the small energy spacing regime. We shall show that the challenge can be alleviated by allowing the coupling between the ambient field and the probe to change sign in a suitably controlled way.
Mathematically, the purpose of this paper is to investigate to what extent the assumption of an infinitely long interaction duration is responsible for the disparity between the circular motion effective temperature and the Unruh temperature in the small gap regime. In brief: How long does an observer undergoing circular motion in $2+1$ dimensions need to wait to register approximate thermality when the observer’s internal energy spacing is small? Physically, the answer has clear resource implications for any experiment.
@@
Switching functions with sign changes appear not to have received attention in the literature. However, there is nothing that prevents us a priori from considering such switching functions, or from constructing them in laboratory experiments. In fact, they arise naturally in entanglement harvesting protocols in which causally disconnected local laser pulses become entangled through their interaction with, for example, an electromagnetic field [48] or a Bose-Einstein condensate [34]. In this setting, the switching function is given by the coherent amplitude of the electric field of the laser, which can be given any profile, and in particular, it can be engineered to change sign. In the Bose-Einstein condensate experiment proposed in [32, 33], the engineering would be related to the detuning of the probe laser beam sidebands around a resonance frequency of the condensate.
@@ 2.2 Stationary response and the long time limit
Consider now the limit in which the interaction operates for a long time with approximately constant coupling strength. As $\chi$ has by assumption compact support, we may formalise the long time limit by considering a family of switching functions $\chi_{\lambda}$, where $\lambda$ is a positive parameter, such that the support of $\chi_{\lambda}$ is an interval with length proportional to $\lambda$ as $\lambda\to\infty$, and $\chi_{\lambda}$ is approximately constant within this interval as $\lambda\to\infty$, at least in some averaged sense. Specific implementations of this limit were introduced in [26], under the assumption that $\chi_{\lambda}$ is non-negative, and we will employ these implementations in Sections 4 and 6. These implementations satisfy
@@ 2.3 Effective temperature
For uniform linear acceleration with proper acceleration $a$, with the field initially prepared in the Minkowski vacuum, $T(E)$ is independent of the energy $E$ and equal to the Unruh temperature $T_{U}=a/(2\pi)$ [68, 22]. The state $\rho(E)$ (2.8) in this case is thus a genuine thermal Gibbs state in the temperature $T_{U}$, and (2.9) is the detailed balance relation between excitation and de-excitation rates in thermal equilibrium [25, 67]. For other types of non-inertial stationary motion, $\rho(E)$ (2.8) is not a thermal state in the Gibbs sense since $T(E)$ depends on $E$; however, for sufficiently narrow intervals in $E$, we may view $\rho(E)$ as an approximate Gibbs state in the energy-dependent temperature $T(E)$. We refer to $T(E)$ as the (infinite time) detailed balance temperature. Studies of $T(E)$ in a range of situations are given in [31, 5, 39, 12, 10, 54, 11].
@@ 3 Asymptotically scaled switching families
using Plancherel’s theorem. The particular case (3.3) follows immediately with $V(E,\omega)=\widehat{\mathcal{W}}(E+\omega)$, and the last statement (3.4) is just the fact that the existence of a double limit implies the existence of the limit in $\lambda$ with $E$ held fixed. ∎
@@ 8 Conclusions
In the simultaneous limit of small gap and long interaction time, most of our examples assumed that the gap and the interaction time are related by an inverse power law, and the examples recovering a finite small gap temperature had power-laws where the interaction time goes to infinity more slowly than the inverse gap. We discussed in Section 7.2.1 and Appendix C an adiabatic scaling example where the interaction profile is stretched more slowly than any inverse power of the gap, but we also noted that in such examples the interaction profile itself is likely to have a weak falloff at early and late times, and hence to be of less interest for analogue spacetime experiments. We leave further discussion of non-power-law scalings to future work.
@@
We considered a massless scalar field in $(2+1)$-dimensional Minkowski spacetime because this field appears in the experimental proposals that employ a planar Bose-Einstein condensate or a superfluid Helium surface [56, 32, 5, 71, 9, 33]. A massless fermion field in $(2+1)$-dimensional Minkowski spacetime can be simulated in an optical lattice, and this system has been proposed as an analogue spacetime simulation of the linear motion Unruh effect [57, 42]. Were it possible to use the optical lattice for a simulation of the circular motion Unruh effect, such that the signal extraction can be modelled by an UDW detector coupled linearly to the scalar density of the spinor field, the detector analysis would be as for a massless scalar field in $5+1$ dimensions [50], and there would be no suppression of the signal in the small gap and long time regime [54]. A Maxwell field in $(2+1)$-dimensional Minkowski spacetime is not known to us to arise in an analogue spacetime simulation, but were such a system found, and proposed for simulating the circular motion Unruh effect, such that the signal extraction is modelled by an UDW detector with a dipole interaction, the linear acceleration results in Section 9.3 of [66] suggest that the outcomes would be similar to a massless scalar field in $4+1$ dimensions, and hence display no suppression of the signal in the small gap and long time regime [54].
@@ Acknowledgements
We thank Benito Juárez-Aubry, Cameron Bunney, Cisco Gooding, Robert Mann, Rick Perche, Bill Unruh and all the members of the Nottingham Gravity Laboratory for stimulating discussions throughout the course of this work. We thank the anonymous referees for helpful comments. DV-C thanks CONACyT and the WW Smith Trust for financial support. The work of CJF was supported by the Engineering and Physical Sciences Research Council [grant number EP/Y000099/1]. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. The authors have benefited from the activities of COST Action CA23115: Relativistic Quantum Information, funded by COST (European Cooperation in Science and Technology). For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.
@@ B.1 $\alpha>1$
Suppose $\alpha>1$. Then $\eta(|E|)\to\infty$ as $E\to 0$, and from (B.3) we have $|E|/\eta(|E|)\to 0$ as $E\to 0$. Using the evenness of $\bigl|\widehat{\psi}\bigr|^{2}$, we may write (B.1) as
@@ B.3 $\alpha<1$
$\mathcal{B}(E)=\eta(|E|)\int_{-1}^{1}\mathrm{d}v\,\frac{1-\cos\big(\eta(|E|)v\big.)}{\eta^{2}(|E|)v^{2}}|\widehat{\psi}(|E|v)|^{2},$ (B.10)
@@
$\frac{1}{|E|}\int_{-\eta(E)}^{\eta(E)}\mathrm{d}u\,{|\widehat{\chi}(u)|}^{2}$ $=\frac{2S}{|E|}{\textstyle{E_{1}\!\left(\log\bigl(\frac{S}{\eta(E)}\bigr.)\right)}}\bigl(1+o(1)\bigr)$
$=\frac{2S}{|E|}\!\left(\frac{\frac{\eta(E)}{S}}{\log\bigl(\frac{S}{\eta(E)}\bigr.)}\right)\bigl(1+o(1)\bigr)$
$=\frac{2\lambda(E)}{\log(S/|E|)-\log\bigl(\lambda(E)\bigr.)}\bigl(1+o(1)\bigr),$ (C.4)
@@
Finally, combining (D.3a) and (D) gives $I(v)=-\frac{2\sqrt{3}}{\pi^{2}}\log\bigl(1-v\bigr.)+O(1)$ as $v\to 1$.
@@
$J(\gamma)=\sqrt{3}\log\bigl(\tfrac{2}{\sqrt{3}}\gamma\bigr.)+\int_{0}^{\infty}\mathrm{d}z\left(\frac{1}{\sqrt{1-\operatorname{\text{sinc}}^{2}\!z}}-\frac{\sqrt{3}}{z(1+z)}-1\right)+o(1),$ (D.9)
@@ E.2 $\chi_{\lambda,\delta}$ is an ASSF
[…]
@@ E.3 $\chi_{\lambda,\delta}$ satisfies SFS when $0<\alpha<2/3$ and $\delta>3/2$
$=\int_{{\mathbb{R}}^{2}}\mathrm{d}u\,\mathrm{d}v\,\cos\bigl(\omega(u-v)\bigr.)\chi(u/\lambda)\chi(v/\lambda)f(u/\lambda^{\delta})f(v/\lambda^{\delta})$
$=\lambda^{2}\int_{{\mathbb{R}}^{2}}\mathrm{d}x\,\mathrm{d}y\,\cos\bigl(\omega\lambda(x-y)\bigr.)\chi(x)\chi(y)f(\lambda^{1-\delta}x)f(\lambda^{1-\delta}y)\,$, (E.6)
[…]